Sophieee:

Expand the squares of the following trinomials. (i) a + 2b + 3c (ii) 2x + 3y + 4z (iii) x + 2y – 3z (iv) 3a – 4b – c (v) 1 – x - 1x (vi) 1 – a – a2

surjithayer:

\[(x+y )^2=x^2+2xy+y^2,(x+y+z)^2=x^2+y^2+z^2+2xy+2yz+2zx\]

surjithayer:

i will solve (vi) \[(1-a-a^2)^2=1^2+(-a)^2+(-a^2)^2+2(1)(-a)+2(-a)(-a^2)+2(-a^2)(1)\] \[=1+a^2+a^4-2a+2a^3-2a^2\] \[=a^4+2a^3-a^2-2a+1\]